# Fast nonparametric feature selection with error control using integrated path stability selection

### v1

@@ Acknowledgements
in Equation S1.4 is equal to a fixed cutoff $C$. As noted in the main text, we always use $C=0.05$, but results are largely independent of this choice (Figures S1 and S2). Figures S3 and S4 show IPSS is also indifferent to $\mu$. Here, the parameter $\delta$ determines the measure $\mu_{\delta}(d\lambda)=z_{\delta}^{-1}\lambda^{-\delta}d\lambda$, where the normalizing constant $z_{\delta}$ is easily computed in closed form (Melikechi and Miller, 2024). The probability measure that we always use in the main text, $\mu_{1}$, averages over $\Lambda$ on a log scale, while $\mu_{0}$ averages on a linear scale. The insignificance of $C$ and $\mu$ is unsurprising: Intuitively, the efp scores for IPSS depend primarily on $f$ and the integrand in $\mathcal{I}(\Lambda)$. The actual value of $\mathcal{I}(\Lambda)$ is much less important since the efp scores depend on the relative quantities $\mathcal{I}(\Lambda)/\int_{\Lambda}f(\hat{\pi}_{j}(\lambda))\mu(d\lambda)$ rather than on $\mathcal{I}(\Lambda)$ itself. Since $C$ and $\mu$ only affect the value of the bound, not the integrand, they contribute little to the actual performance of IPSS.
Figure S1. Dependence on cutoff (regression). IPSSGB is applied to regression data simulated as in Section 4 for cutoff values $C\in\{0.025,0.05,0.075,0.1,0.15\}$, averaged over 100 trials. There is little difference between the average true and false positive values for different choices of $C$ over a wide range of target E(FP) values. Figures S2, S3, and S4 show similar results for classification and dependence on $\mu$.
Figure S2. Dependence on cutoff (classification). IPSSGB is applied to classification data simulated as in Section 4 for cutoff values $C\in\{0.025,0.05,0.075,0.1,0.15\}$, averaged over 100 trials. There is little difference between the average true and false positive values for different choices of $C$ over a wide range of target E(FP) values.
Figure S3. Dependence on $\mu$ (regression). IPSSGB is applied to regression data simulated as in Section 4 for cutoff values $\delta\in\{0,0.5,1,1.5,2\}$, averaged over 100 trials, where $\delta$ determines to the measure $\mu_{\delta}(d\lambda)=z_{\delta}^{-1}\lambda^{-\delta}d\lambda$. There is little difference between the average true and false positive values for different choices of $\delta$ over a wide range of target E(FP) values.
Figure S4. Dependence on $\mu$ (classification). IPSSGB is applied to classification data simulated as in Section 4 for cutoff values $\delta\in\{0,0.5,1,1.5,2\}$, averaged over 100 trials, where $\delta$ determines to the measure $\mu_{\delta}(d\lambda)=z_{\delta}^{-1}\lambda^{-\delta}d\lambda$. There is little difference between the average true and false positive values for different choices of $\delta$ over a wide range of target E(FP) values.
S1.4.  Preselection.
Many feature selection algorithms, including most in Section 3, employ some form of screening, or preselection, as an initial step in the selection process. This can be especially helpful in high dimensions for reducing noise and runtime. When $p>200$, we preselect features for IPSSGB, IPSSRF, and the other stability selection-based methods (IPSSL and SSBoost, detailed below) by running their respective selection algorithms on the full dataset three times—that is, computing $\Phi_{Z_{1:n}}$ three times—and keeping only the $p_{\text{pre}}=\max\{200,p/20\}$ features with the largest average scores across all three trials. Note that $\Phi_{Z_{1:n}}$ is stochastic, given $Z_{1:n}$, due to randomness in the gradient boosting and random forest algorithms. This preselection step does not affect the theoretical control on E(FP) since $\lvert\hat{S}_{\mathrm{IPSS},\mathrm{pre}}\cap S^{c}\rvert=\lvert\hat{S}_{\mathrm{IPSS},\mathrm{pre}}\cap S^{c}_{\mathrm{pre}}\rvert$, where $\hat{S}_{\mathrm{IPSS},\mathrm{pre}}$ are the features selected by IPSS using only the preselected features, and $S^{c}_{\mathrm{pre}}$ are the preselected features in $S^{c}$. Of course, preselection risks throwing away important features, potentially increasing the number of false negatives. However, we have yet to observe a case in which an important feature that was not in the top $p_{\mathrm{pre}}$ preselected features was selected by IPSS when the full set of features was used instead. In other words, a feature that does not make the top $p_{\mathrm{pre}}$ is highly unlikely to be selected by IPSS regardless of whether or not preselection is used.
S2.  Simulation details and additional results
@@
Algorithm S2 describes the procedure we use to generate data in the simulation study in Section 4. See also Figure 1 in the main text for a diagrammatic representation of this process. In all steps, “randomly select” means select a parameter uniformly at random from its domain, which are shown in Table S1. The randomized function $f_{\theta}:\mathbb{R}\to[-1,1]$ that links the features to the response is
@@
Parameter $\alpha$ $\beta$ $\gamma$ $\delta_{1}$ $\delta_{2}$ $p_{\mathrm{true}}$ SNR $u$
Range $(0.5,1.5)$ $(-1,1)$ $(1,3)$$\{-1,1\}$$\{-1,1\}$$\{10,\dots,30\}$ $(0.5,2)$ $(1,3)$
@@
1:RNA-seq data $X_{\mathrm{full}}\in\mathbb{R}^{596\times 6426}$, number of samples $n$, number of features $p$, number of true features $p_{\mathrm{true}}$, signal-to-noise ratio SNR, function parameter domain $\Theta$.
@@
4:Randomly select true features $S\subseteq\{1,\dots,p\}$ with $\lvert S\rvert=p_{\mathrm{true}}$.
5:Randomly select $G\in\{\lfloor p_{\mathrm{true}}/2\rfloor,\dots,p_{\mathrm{true}}\}$. Partition $S$ into $G$ disjoint groups, $S=\bigsqcup_{g=1}^{G}S_{g}$.
@@
Figure S5. Regression simulation. Empirical E(FP) and E(TP) versus target E(FP) for each IPSS method. The dotted black line shows where E(FP) = target E(FP).
Figure S6. Classification simulation. Empirical E(FP) and E(TP) versus target E(FP) for each IPSS method. The dotted black line shows E(FP) = target E(FP).
@@
One benefit of IPSS is that changing the target E(FP) does not require rerunning the algorithm. Figures S5 and S6 show E(TP) and E(FP) as functions of the target E(FP). IPSSGB has the best performance in terms of true positives across essentially all target E(FP) values. IPSSGB and IPSSRF perform well in the regression setting, and both slightly overshoot the target E(FP) in classification for values less than 6. IPSSL performs poorly in general, having relatively few true positives on average and false positives that exceed the target E(FP).
S3.  Additional ovarian cancer results
In Section S3.1, we expand upon the ovarian cancer results presented in Section 5 of the main text, namely identifying miRNAs related to prognosis, and miRNAs related to miR-150, respectively. In Section S3.2, we perform a similar analysis but for RNA-seq data from the same cohort, identifying genes related to ovarian cancer prognosis, and genes related to one of the genes, AKT2, that was found to be related to prognosis.
S3.1.  MicroRNA and ovarian cancer.
@@
We briefly summarize results from the literature that link the miRNAs in Table 2 to ovarian cancer, with a specific focus on prognosis. For many of these miRNAs, there are numerous studies connecting them to ovarian cancer beyond what we present. Thus, to roughly quantify the relevance of the miRNAs in Table 2 to ovarian cancer, we searched the specific miRNA together with “ovarian cancer” in the Web of Science database, and report both the number of publications and the total number of citations across all returned results. We emphasize that these metrics are less important than the papers themselves. For example, citation counts favor older publications, and our search criterion does not guarantee that the returned results are relevant to the problem at hand (though the summaries below show that at least some of them are).
Table S2 is similar to Table 2, except that it shows the total number of publications and citations from our Web of Science literature search and includes some $q$-values that were omitted from Table 2, since Table 2 only reported the five miRNA with the lowest $q$-values for each method. Dashed lines indicate that the feature was not identified by the corresponding method at all.
miRNA Citations Publications IPSSGB IPSSRF IPSSL
miR-150 1635 26 0.28 0.41 0.2
miR-93 1270 28 0.36
miR-30d 596 12 0.43 0.36 0.16
miR-96 338 10 0.2 0.53 0.09
miR-342 201 9 0.36 0.09
miR-1270 175 5 0.43 0.27 0.08
miR-934 27 2 0.12
miR-1301 11 3 0.46
miR-1-2 5 1 0.07
Table S2. MicroRNA and prognosis. MiRNA are ordered by citations. A missing $q$-value means that miRNA was not selected by the corresponding version of IPSS.
@@
miR-96. Liu et al. (2019) found that overexpression of miR-96 promotes cell proliferation and migration in ovarian cancer cells. They conclude that targeting miR-96 is a potentially promising strategy for treating ovarian cancer. They also report that miR-96 inhibits phosphorylation of AKT, a gene identified by IPSSGB as being relevant to ovarian cancer prognosis (Section S3.2). Yang et al. (2020) found that individuals with low-levels of miR-96 “suffered more advanced tumor staging and a worse overall survival” and also identified miR-96 as a potential therapeutic target.
@@
miR-934. Hu et al. (2019) found that miR-934 promotes ovarian cancer cell proliferation and inhibits the death of ovarian cancer cells by targeting the gene BRMS1L, making it is a potential biomarker or therapeutic target for ovarian cancer.

### v2

@@ Acknowledgements
in Equation S1.4 is equal to a fixed cutoff $C$. As noted in the main text, we always use $C=0.05$, but results are largely independent of this choice (Figures S12 and S13). Figures S14 and S15 show IPSS is also robust to the parameter $\delta$ that determines the measure $\mu_{\delta}(d\lambda)=z_{\delta}^{-1}\lambda^{-\delta}d\lambda$, where the normalizing constant $z_{\delta}$ is easily computed in closed form (Melikechi and Miller, 2024). The probability measure $\mu_{1}$ averages over $\Lambda$ on a log scale, while $\mu_{0}$ averages on a linear scale.
The insignificance of $C$ and $\delta$ is unsurprising: Intuitively, the efp scores for IPSS depend primarily on $f$ and the integrand in $\mathcal{I}(\Lambda)$. The actual value of $\mathcal{I}(\Lambda)$ is much less important since the efp scores depend on the relative quantities $\mathcal{I}(\Lambda)/\int_{\Lambda}f(\hat{\pi}_{j}(\lambda))\mu(d\lambda)$ rather than on $\mathcal{I}(\Lambda)$ itself. Since $C$ and $\mu$ only affect the value of the bound, not the integrand, they contribute little to the actual performance of IPSS, as indicated by the sensitivity analyses in Section S5.
Method $B$ $C$ $f(x)$ $\delta_{\text{reg}}$ $\delta_{\text{class}}$
IPSSGB $100$ $0.05$ $(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$ $1.25$ $1$
IPSSRF $50$ $0.05$ $(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$ $1.25$ $1.25$
Table S1. Default IPSS parameters. Both IPSSGB and IPSSRF always use $C=0.05$ and $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$. IPSSRF uses $B=50$ to reduce runtimes without any noticeable difference in selection performance. The parameters $\delta_{\text{reg}}$ and $\delta_{\text{class}}$ determine the measure $\mu_{\delta}$ in regression and classification problems, respectively.
S2.  Overview of methods and implementation details
We describe the preselection procedure used to improve feature selection (Section S2.1) and provide implementation details for each feature selection method considered in this work (Section S2.2).
S2.1  Preselection.
Many feature selection methods employ some form of screening, or preselection, as an initial step in the selection process. This can be especially helpful in high dimensions for increasing power and reducing runtimes. For many of the methods in Section S2.2, we preselect features by running a randomized importance function on the full dataset three times—that is, computing $\Phi_{Z_{1:n}}$ three times—and keeping only the $p_{\text{pre}}$ features with the largest average scores across all three trials. For example, preselection for IPSSRF entails fitting three random forests to the full dataset and keeping the $p_{\text{pre}}$ features with the largest average importance scores.
For IPSS, this preselection step does not affect the theoretical control on E(FP) since $\lvert\hat{S}_{\mathrm{IPSS},\mathrm{pre}}\cap S^{c}\rvert=\lvert\hat{S}_{\mathrm{IPSS},\mathrm{pre}}\cap S_{\mathrm{pre}}^{c}\rvert$, where $\hat{S}_{\mathrm{IPSS},\mathrm{pre}}$ are the features selected by IPSS using only the preselected features, and $S_{\mathrm{pre}}^{c}$ are the preselected features in $S^{c}$. Of course, preselection risks discarding important features, potentially increasing the number of false negatives. In practice, however, we find that preselection actually helps IPSS, stability selection, and model-X knockoffs identify more true positives while still controlling false discoveries. This is because preselection gets rid of many noisy features, making it easier for these methods to detect the true signal.
In Section S2.2, we describe the preselection parameters for each method, which were determined by extensive testing on simulated data. We do not apply preselection when implementing Boruta, RFEGB, RFHT, Vita, and VSURF since these methods include their own internal screening steps.
S2.2  Methods.
We describe how each method in Table S2 is implemented in this work.
Method Package Error control Base method Non-default settings
IPSSGB ipss (Python) Boosting
IPSSRF ipss (Python) Random forest
IPSSL1 ipss (Python) Lasso
KOGLM knockoff (R) GLM
KOL1 knockoff (R) Lasso
KORF knockoff (R) Random forest
DeepPINK knockpy (Python) Neural network
SSBoost XGBoost (Python) Boosting $\texttt{assumption}=r\mathrm{-concave}$
with stabs (R) $\tau=0.75$
KOBT KOBT (R) Boosting $\texttt{num}=100$, $\texttt{bound}=200$,
[…]
RFHT rfvimptest (R) Random forest
Boruta Boruta (R) Random forest
RFEGB scikit-learn (Python) Boosting
Vita vita (R) Random forest $p$-value threshold $=0$
VSURF VSURF (R) Random forest VSURF_pred
Table S2. Feature selection methods. Software packages are listed with the language used to implement them in parentheses. Details about each method, including descriptions of their non-default settings, are in Sections S2.2.1S2.2.7. For methods with no non-default settings, we use the default settings in their respective packages.
S2.2.1  IPSSGB.
The IPSS-related parameters used to implement IPSSGB are in Table S1. For preselection, we use gradient boosting as the baseline selection algorithm and set $p_{\text{pre}}=100$. We implement gradient boosting using XGBoost (Chen and Guestrin, 2016). All XGBoost parameters are set to their default values except for two changes: The proportion of features considered when splitting each node (called colsample_bynode in XGBoost and often $\mathtt{mtry}$ elsewhere) is changed from $1$ to $1/3$, and the maximum depth of each tree (max_depth) is changed from $6$ to $1$, making each tree a stump. The latter change significantly improved the performance of IPSSGB, both in terms of speed and feature selection.
S2.2.2  IPSSRF.
The IPSS-related parameters used to implement IPSSRF are in Table S1. For preselection, we use random forests as the baseline selection algorithm and set $p_{\text{pre}}=100$. We implement random forests using scikit-learn (Pedregosa et al., 2011). All random forest parameters are set to their default values except for two changes: The proportion of features considered when splitting each node (called max_features in scikit-learn) is changed from $1$ to $1/10$, and the number of trees used to build each random forest (n_estimators) is changed from $100$ to $50$. These changes improved the efficiency of IPSSRF without sacrificing its feature selection performance.
S2.2.3  IPSSL1.
As discussed in the main text, IPSSL1 is a parametric version of IPSS based on $\ell^{1}$-regularization (Melikechi and Miller, 2024). For regression, the baseline algorithm is lasso (Tibshirani, 1996), and for classification, it is $\ell^{1}$-regularized logistic regression (Friedman et al., 2010). All parameters are set to their default values in the ipss Python package: https://pypi.org/project/ipss/.
S2.2.4  Model-X knockoffs (KOGLM, KOL1, KORF, DeepPINK, KOBT).
Model-X knockoffs work by first constructing knockoffs $\tilde{X}=(\tilde{X}_{1},\dots,\tilde{X}_{p})$ of the original features $X=(X_{1},\dots,X_{p})$. By definition, $\tilde{X}$ must satisfy (i) the joint distribution of $(X,\tilde{X})$ is invariant under pairwise exchanges of the original features $X_{j}$ and their corresponding knockoffs $\tilde{X}_{j}$, and (ii) $\tilde{X}$ is conditionally independent of the $Y$ given $X$ (see Definition 2 in Candes et al. (2018)). Once knockoffs are constructed, feature importance scores—called feature statistics in the knockoffs literature—are computed for all of the original and knockoff features and subsequently used to select original features in a way that controls the FDR. Like IPSS, any feature importance function can be used.
As noted in the main text, model-X knockoffs requires knowledge of the joint distribution of $X$. In our multivariate Gaussian simulations in Section 3, all of the model-X knockoffs methods are implemented using the true, known joint distribution. In the remaining examples, where the joint distribution of $X$ is not known, we use the default methods for constructing approximate model-X knockoffs (for example, second-order Gaussian knockoffs in the knockoff R package).
KOGLM, KOL1, and KORF are all implemented using the R package knockoff. KOGLM uses feature importance scores from a generalized linear model (GLM); we find it performs best when using random forests for preselection with $p_{\text{pre}}=200$. KOL1 uses feature importance scores from lasso (regression) or $\ell^{1}$-regularized logistic regression (classification). Like IPSSL1, we use lasso (or logistic regression) for preselection, with $p_{\text{pre}}=200$. KORF uses feature importance scores from random forests; like KOGLM, we find it performs best when using random forests for preselection with $p_{\text{pre}}=200$.
DeepPINK uses deep neural networks to construct feature importance scores. We implement it using the Python package knockopy and random forests for preselection with $p_{\text{pre}}=100$.
KOBT is implemented with the R package KOBT. It uses boosted tree models to construct importance scores. We tested KOBT numerous times on simulated data with many different tuning and preselection parameters (including no preselection), but found that KOBT consistently and dramatically exceeded its target FDR. For this reason, we omit its results from all plots and tables in this work.
Finally, we tested all of the model-X knockoffs methods without using preselection. In these cases, power was often extremely low and the runtimes were much longer.
S2.2.5  SSBoost.
The closest method to IPSSGB in terms of its underlying approach is that of Hofner et al. (2015), referred to here as SSBoost. Unlike IPSSGB—which uses importance scores from gradient boosting—SSBoost applies stability selection to choose the number of features used per boosting run. Furthermore, IPSSGB uses IPSS to construct efp scores, whereas SSBoost uses a version of stability selection introduced by Shah and Samworth (2013). This is perhaps the most significant difference since the efp scores for IPSS have much tighter bounds than those for other forms of stability selection (Melikechi and Miller, 2024). From a practical standpoint, this causes other versions of stability selection to identify fewer important features than IPSS.
Hofner et al. (2015) provide code for SSBoost that combines the R packages mboost (Hofner et al., 2014) and stabs (Hofner and Hothorn, 2017) in the form of a worked example, but the mboost implementation of boosting was prohibitively slow in the dimensions we consider. By adapting their code to use XGBoost in place of mboost for boosting and by preselecting features using gradient boosting with $p_{\text{pre}}=150$, we were able to reduce SSBoost runtimes considerably with no apparent change in results. The XGBoost parameters used for SSBoost are the same as those used for IPSSGB (Section S2.2.1). For the stability selection part of SSBoost, we use the default parameters in stabs, and the selection threshold is set to $\tau=0.75$, which is the middle of the interval $(0.6,0.9)$ recommended by Meinshausen and Bühlmann (2010).
S2.2.6  RFHT.
We tested RFHT (Coleman et al., 2022), which achieves theoretical error control by using random forests for hypothesis testing. However, one test run with default parameters on simulated data with 500 samples, 500 features, and 20 true features took over 51 minutes, returning 15 true positives and 43 false positives. By contrast, IPSSGB with a target FDR of $0.2$ took 11 seconds and returned 10 true positives and 0 false positives on the same data. This method is omitted from our studies because its performance does not appear to justify its excessive runtime.
S2.2.7  Methods without false discovery control (Boruta, RFEGB, Vita, VSURF).
Boruta, Vita, and VSURF are implemented using R packages of the same names. Boruta is run with default parameters. For Vita, we set the $p$-value threshold to 00. For VSURF, we use the function VSURF_pred rather than VSURF_interp to select the final set of features (Genuer et al., 2010). Both of these choices favor sparsity, which aligns well with our simulation designs. As noted in the main text, VSURF is too computationally expensive to include in our $p=2000$ and $5000$ simulation studies; see also Table S3.
We implement RFEGB by combining XGBoost and scikit-learn. On simulated Gaussian data with $n=250$ and $p=500$, one run of RFEGB took over 12 minutes when removing five features per iteration (the default is one feature removed per iteration, which takes approximately 5 times as long). For comparison, IPSSGB ran in 5 seconds on the same data and had far fewer false positives and similar power. Due to its high computational cost, poor performance, and many tuning parameters (which is true of recursive feature elimination in general), RFEGB is largely omitted from this work.
S3.  Simulation results and details
We present additional simulation results (Section S3.1) and RNA-seq simulation details (Section S3.2).
S3.1  Additional simulation results.
Figure S1 shows the $n=250$ multivariate Gaussian simulation results, described in Section 3.2. Figure S2 shows the $p=500$, $2000$, and $5000$ RNA-seq simulation results for classification, described in Section 3.3. Tables S3 and S4 show the average runtimes of each method in each simulation experiment.
Figure S1. Gaussian simulation results ($n=250$). See Figure 1 for details.
Figure S2. RNA-seq simulation results (classification). See Figure 2 for details.
Method MG($n=250$) MG($n=500$) RNA($p=500$) RNA($p=2000$) RNA($p=5000$)
IPSSGB 4.4 (0.11) 6.3 (0.01) 6.7 (0.23) 9.0 (0.43) 14.6 (0.45)
IPSSRF 4.5 (0.05) 7.5 (0.10) 7.2 (0.14) 10.8 (0.29) 19.6 (0.63)
IPSSL1 1.6 (0.23) 1.0 (0.05) 4.0 (0.91) 3.1 (0.66) 6.4 (0.76)
KOGLM 10.8 (0.92) 12.4 (0.91) 10.0 (1.92) 12.6 (1.50) 19.2 (1.66)
KOL1 11.2 (0.88) 12.0 (1.04) 7.5 (1.27) 6.7 (0.93) 7.2 (1.16)
KORF 12.8 (1.16) 14.3 (1.01) 4.2 (0.50) 7.4 (0.22) 13.9 (0.50)
DeepPINK 5.0 (3.55) 6.0 (2.79) 5.9 (4.17) 8.8 (0.21) 15.4 (0.42)
SSBoost 4.1 (0.04) 6.7 (0.03) 6.7 (0.04) 8.1 (0.15) 13.7 (0.08)
Boruta 42.2 (3.09) 120.7 (7.93) 42.0 (5.96) 57.0 (6.51) 84.5 (6.66)
Vita 9.5 (0.19) 24.2 (0.40) 20.6 (0.47) 83.5 (4.09) 195.3 (4.34)
VSURF 196.8 (8.16) 548.6 (23.68) 286.7 (72.36)
RFHT 3087*
RFEGB 749*
Table S3. Average runtimes (in seconds) over 100 trials for each regression experiment. MG stands for Multivariate Gaussian. Standard deviations are in parentheses. Recall that $p=500$ in the MG experiments and $n=500$ in the RNA experiments, and that VSURF was too computationally expensive to include when $p=2000$ and $5000$As discussed in Sections S2.2.6 and S2.2.7, RFHT and RFEGB are largely omitted due to their poor selection performance and excessive runtimes in initial tests (shown in the table). Hence, their remaining entries are blank.
Method MG($n=250$) MG($n=500$) RNA($p=500$) RNA($p=2000$) RNA($p=5000$)
IPSSGB 4.3 (0.04) 6.5 (0.11) 6.6 (0.07) 7.5 (0.20) 14.3 (0.42)
IPSSRF 4.2 (0.02) 7.0 (0.05) 6.8 (0.10) 9.9 (0.22) 17.7 (0.49)
IPSSL1 5.9 (0.15) 6.0 (0.22) 8.4 (0.86) 9.3 (1.17) 13.1 (1.25)
KOGLM 10.6 (0.10) 12.0 (0.14) 9.3 (1.57) 11.8 (1.67) 17.4 (1.64)
KOL1 10.7 (0.12) 11.9 (0.21) 7.2 (1.13) 6.0 (0.83) 6.3 (0.80)
KORF 10.1 (0.95) 12.1 (0.17) 3.1 (0.10) 5.9 (0.18) 11.3 (0.38)
DeepPINK 4.7 (0.29) 5.7 (0.28) 5.2 (0.10) 8.0 (0.24) 13.3 (0.43)
SSBoost 4.0 (0.01) 6.6 (0.04) 6.7 (0.02) 8.2 (0.04) 13.8 (0.11)
Boruta 27.8 (1.81) 76.6 (6.29) 25.4 (3.91) 37.4 (4.51) 58.6 (3.99)
Vita 4.8 (0.06) 10.7 (0.06) 9.2 (0.16) 35.3 (0.72) 89.1 (4.59)
VSURF 78.1 (4.14) 190.9 (8.93) 77.0 (28.08)
Table S4. Average runtimes (in seconds) over 100 trials for each classification experiment. MG stands for Multivariate Gaussian. Standard deviations are in parentheses. Recall that $p=500$ in the MG experiments and $n=500$ in the RNA experiments. VSURF was too computationally expensive to include when $p=2000$ and $5000$.
S3.2  RNA-seq simulation details.
@@
Algorithm S2 describes the data generating procedure for the RNA-seq simulation studies (Section 3.3), which is also depicted in Figure S3. In all steps, “randomly select” means select a parameter uniformly at random from its domain, which are shown in Table S5. The randomized function $f_{\theta}:\mathbb{R}\to[-1,1]$ that links the features to the response is
@@
Parameter $\alpha$ $\beta$ $\gamma$ $\delta_{1}$ $\delta_{2}$
Range $(0.5,1.5)$ $(-1,1)$ $(1,3)$$\{-1,1\}$$\{-1,1\}$
@@
1:RNA-seq data $X_{\mathrm{full}}\in\mathbb{R}^{596\times 6426}$, number of samples $n$, number of features $p$, number of true features $\lvert S\rvert$, signal-to-noise ratio SNR, function parameter domain $\Theta$.
@@
4:Randomly select $\lvert S\rvert$ true features $S\subseteq\{1,\dots,p\}$.
5:Randomly select $G\in\{\lfloor\lvert S\rvert/2\rfloor,\dots,\lvert S\rvert\}$. Partition $S$ into $G$ disjoint groups, $S=\bigsqcup_{g=1}^{G}S_{g}$.
@@
[…]
Figure S3. Simulation diagram. Rows and columns are randomly selected from the full RNA-seq dataset to create a matrix $X$ whose columns are split into important features, $X_{S}$, and unimportant features, $X_{S^{c}}$. The columns of $X_{S}$ are further partitioned into $G$ groups; the above figure shows $G=3$. The features in each group are summed to obtain $\xi_{g}$, and a different realization $f_{\theta_{g}}$ of $f_{\theta}$ is applied to $\xi_{g}$ for each $g$. For regression, response is the sum of the group-specific signals, $f_{\theta_{g}}(\xi_{g})$, plus noise.
@@
Figure S5. Distributions of five randomly selected genes from the RNA-seq dataset. Features in the ovarian cancer RNA-seq dataset follow a variety of empirical marginal distributions. For example, the standardized empirical distributions of the five randomly selected genes above, from left to right, are relatively flat, skewed left, approximately Gaussian, skewed right, and contain outliers. Furthermore, the genes exhibit complex correlation structures, with maximum and average absolute pairwise correlations of approximately 0.95 and 0.17 after standardization, respectively.
S4.  Cancer study results
We describe our ovarian cancer and glioma studies (Sections S4.1 and S4.2) and present our literature search and cross-validation results (Sections S4.3 and S4.4).
S4.1  Ovarian cancer.
For the ovarian cancer cohort, we study the following feature and response combinations (the number of samples and features are shown in parentheses): (i) miRNAs and prognosis ($n=442$, $p=585$), (ii) miRNAs and tumor purity ($n=451$, $p=585$), (iii) miRNAs and miR-150 ($n=453$, $p=584$), (iv) genes and prognosis ($n=549$, $p=6426$), and (v) genes and AKT2 ($n=569$, $p=6425$). In all cases, missing values were removed. Gene expression levels are measured by RNA-seq. We chose miR-150 and the gene AKT2 as responses in studies (iii) and (v), respectively, because literature searches indicated that both are highly related to ovarian cancer. Literature search results are only reported for clinical responses (prognosis and tumor purity).
S4.2  Glioma.
For the glioma cohort, we study the following feature and response combinations (the number of samples and features in each dataset are shown in parentheses): (i) miRNAs and prognosis ($n=477$, $p=787$), (ii) miRNAs and miR-155 ($n=512$, $p=786$), (iii) genes and prognosis ($n=625$, $p=10,058$), and (iv) genes and FOXM1 ($n=669$, $p=10,057$). In all cases, missing values were removed. Gene expression levels are measured by RNA-seq. The original glioma RNA-seq dataset contains over $20,000$ genes; we only study the roughly $10,000$ genes in the top 50th percentile of average gene expression (discarding lowly expressed genes is common practice). We chose miR-155 and the gene FOXM1 as responses in studies (ii) and (iv), respectively, because literature searches indicated that both are highly related to glioma. Literature search results are only reported when the response is prognosis.
S4.3  Literature search.
@@
In each of the forthcoming studies, we perform literature searches to validate our findings. For the microRNA (miRNA) and ovarian cancer prognosis study, we briefly summarize literature supporting each miRNA selected by at least one feature selection method. Providing such summaries for every study is beyond the scope of this work. Thus, to roughly quantify the relevance of selected features in subsequent studies, we searched the feature name, cancer type, and the word “prognosis” in Europe PMC, an open-access database containing millions of life sciences publications (Europe PMC Consortium, 2015). We then report the total number of citations among all articles returned by the search. For example, in Table S6, searching “miR-93” + “ovarian cancer” + “prognosis” in Europe PMC returned 43,996 citations. For each method, we also include the number of features it selected that had above a certain number of citations, below a certain number of citations, and the total number of features it selected below a certain target FDR. We emphasize that these metrics are less important than the papers themselves. Citation counts favor older publications, and our search criterion does not guarantee relevance to the specific problem at hand (though the summaries below suggest that at least some results are meaningful).
Below, we briefly summarize literature relating miRNAs in Table S6 to ovarian cancer. Additional details are available in the associated references.
@@
miR-96. Liu et al. (2019) found that overexpression of miR-96 promotes cell proliferation and migration in ovarian cancer cells. They conclude that targeting miR-96 is a potentially promising strategy for treating ovarian cancer. They also report that miR-96 inhibits phosphorylation of AKT, a gene identified by IPSSGB as being relevant to ovarian cancer prognosis. Yang et al. (2020) found that individuals with low-levels of miR-96 “suffered more advanced tumor staging and a worse overall survival” and also identified miR-96 as a potential therapeutic target.